OpenStudy (afloridagirl):

Vertical angles are congruent to each other! Hint #22 / 6 Set the angle measures equal to one another. \blue{4x + 112^\circ} = \green{9x + 72^\circ}4x+112 ​∘ ​​ =9x+72 ​∘ ​​

sam (.sam.):

Your latex isn't working.

OpenStudy (afloridagirl):

4x+112^degree=9x+72^degree

OpenStudy (afloridagirl):

@.Sam.

OpenStudy (princeharryyy):

4x+112 = 9x+72 112-72=9x-4x 40=5x x=8 deg

OpenStudy (afloridagirl):

thank you\

OpenStudy (princeharryyy):

:)